Hello, I could really use some help on this one. Thanks! Lydia takes ten minutes more to complete an illustration than Tom. The total time taken by both of them is six hours. How much time did each require?
Found 3 solutions by ankor@dixie-net.com, Shin123, MathTherapy:
Answer by ankor@dixie-net.com(22740)

Lydia takes ten minutes more to complete an illustration than Tom. The total time taken by both of them is six hours. How much time did each require?
:
Write a time equation: Let t = Tom's time; Then (t+10) = Lydia's time
:
We will do the problem in minutes, 6 hrs = 360 min:
:
Toms time + Lydia's time = 360 min
t + (t+10) = 360
2t = 360 - 10
2t = 350
t = 350/2
t = 175 min or 2 hr, 55 min
:
Then Lydia would take 185 min or 3 hr 5 min
:
Check 185 + 175 = 360 which is 6 hrs

Answer by Shin123(626)
Answer by MathTherapy(10552)

Hello, I could really use some help on this one. Thanks! Lydia takes ten minutes more to complete an illustration than Tom. The total time taken by both of them is six hours. How much time did each require?
If they worked on the illustration together, and completed it in 6 hours, then the correct answer is: 
If they worked separately on the illustration and total time taken to complete it was 6 hours, then: